A body of mass 20kg was projected into space at an angle of 40° calculate
(a)time taken to reach the maximum height
(b)the maximum height reached
(c)the range
1
Expert's answer
2019-10-10T09:58:05-0400

When body is launched with initial velocity "v_0" at the angle "\theta" to horizon, then


(a) time taken to reach the maximum height


"t=\frac{v_0\sin\theta}{g}"

(b) the maximum height reached


"h_{\rm max}=\frac{v_0^2\sin^2\theta}{2g}"

(c) the range


"R=\frac{v_0^2\sin 2\theta}{g}"